OpenStudy (anonymous):

a) Five times the difference of a number and two is seven more than that number b) Twice a number decreased by four is greater than ten, what numbers satisfy this condition? 2X-4>10 2X>14 X>7 Is b) right? c) Twice a number decreased by two is equivalent to than number increased by five.

OpenStudy (anonymous):

a) 5|a - 2| = a + 7 two solutions: a = 0.5 or a = 4.25

OpenStudy (anonymous):

b) x > 7 c) 2*x - 2 = x + 5 x = 7